OpenStudy (anonymous):

Write the slope-intercept form of the line that passes through the points (7,14) and (9,18).

OpenStudy (anonymous):

to find the slope, use the two points given. slope=m=y2-y1/x2-x1= 18-14/9-7= 4/2=2 now plug everything into the y=mx=b equation y=2x+b (substitute x and y using any of the two coordinates) 14=2(7)+b (isolate for b to find intercept) b= 0 so the equation is y=2x (since b=0)